# CyGym: A Simulation-Based Game-Theoretic Analysis Framework for Cybersecurity

### v1

@@ 3 Cybersecurity Simulator
At the high level, our cybersecurity simulator is comprised of four components: 1) a (hierarchically) networked set of devices, 2) workloads, 3) exploits, and 4) attack mitigation actions. In particular, the analysis is centered around a network comprised of subnets; each subnet, in turn, consists of a network of devices. As such, the key unit of interest in our simulator is whether or not particular devices are compromised at particular points in time. Some devices may be more valuable (for example, servers), and some compromises of these more severe than others (for example, exfiltration of trade secrets compared to a requirement to reboot the device). Generally, we wish to ask questions of the form “how many devices, on average, have been compromised between times $X$ and $Y$?”, “what is the likelihood of particular secrets being exposed to the malicious actor?”, “what is the expected reduction in productivity as a result of cyber-attacks?”, and the like, as we experiment with cyber-attack and defense settings.
@@ 4 Simulation-Based Game-Theoretic Model
We now use the simulation model discussed above to formally define a simulation-based game-theoretic model of cybersecurity encounters. First, there are two players: the defender $\beta$ (e.g., network designer, administrator, etc) and the attacker $\alpha$; these are already explicitly referenced in the simulator described in Section 3). Second, the game takes place over a sequence of discrete time steps $t$ starting at $t=0$ and ending at a predefined horizon $T$. At time 00, the attacker has no knowledge of the network structure $G_{0}$ or device configurations $d_{i}$, and we assume some initial proportion $\eta$ are compromised, all others are clean (i.e., compromise state $c_{i0}$ is a zero vector for all other devices, 1 for the fraction $\eta$). Additionally, we assume that the compromised state $c_{it}$ at any time $t$ after a successful compromise is known to the attacker but not the defender. However, we assume that there is a common knowledge distribution (e.g., stochastic model) of the network generation and device configuration, i.e., $G_{0}\sim\mathcal{D}_{G}$ and $\{d_{i}\}_{i\in D_{0}}\sim\mathcal{D}_{d}$. Consequently, our game entails partial observability of game (device) state by both players.
@@ 4.1 Attacker Actions
[…]
@@ 4.2 Defender Actions
[…]
@@ 5 DOAR: Double Oracle with Actor Response Ascent
[…]
@@
$D^{*}=\{d:a_{d}\neq\mathrm{noop}\},\quad t^{*}=\max_{d}t_{d},\quad e^{*}=\bigcup_{d}e_{d},\quad p^{*}=\text{(from any non-noop }a_{d}).$
@@ 6 Experiments
The environment tracks up to $N_{\max}$ devices, each described by six features—OS code, version number, compromise flag, anomaly score, attacker-knowledge flag, and presence indicator—concatenated (and $-1$-padded) into a full state vector $x\in\mathbb{R}^{6N_{\max}}$. The defender’s observation $o_{d}\in\mathbb{R}^{6N_{\max}}$ masks comprimise information as well as attacker owned and devices not yet added to the network via its evolution procedure. The attacker’s observation $o^{\alpha}\in\mathbb{R}^{4N_{\max}}$ further hides anomaly scores and presence indicators of the defender, drops the attacker-knowledge feature entirely, and masks inactive or unknown devices. The defender’s observations $o^{\beta}$ include anomaly scores and device configurations, but not actual compromise flags.
@@
Defender Clean Clears compromise flag, incurs device busy-time $+0.30$ if compromised, else $-0.01$
Checkpoint Saves state (of all devices) to disk for future rollback $-0.50$
Restore Restores network from last checkpoint (losing intermediate work) $-1.00$
Upgrade Increments application version, incurs busy-time $-1.00$ (device also becomes busy)
Scan Executes ML-based anomaly detection on recent logs $-0.50$
Block / Unblock Toggles network edges to control reachability $-0.50$
Pass No action taken 00
Attacker Attack Deterministically exploits any matching vulnerability on each targeted host $+1.00$ per compromise; $+10.0$ if via domain controller
Probe Discovers previously unknown neighbors via network scanning $+0.10$ per successful discovery
Pass No action taken 00
@@
The network is initialized as a connected Barabási–Albert graph seeded with domain controllers and Windows Server hosts. Initially, a fixed proportion of active nodes are marked compromised at random and a fixed count of nodes are designated attacker-owned; attacker-owned nodes remain permanently compromised and represent devices outside of the network with a connection into the network. They perform no work for the defender. At each time step the environment samples a Poisson$(\lambda)$ number of “events” that either activate or remove devices (while respecting a minimum network size). Newly activated non–attacker nodes are attached via Barabási–Albert preferential attachment; removed nodes simply go offline. Attacker-owned machines remain permanently compromised and continue to influence connectivity. This dynamic topology captures the ebb and flow of hosts in a realistic enterprise network.
@@ 6.1 Results
DOAR $-14.700\pm 0.001$ $-97.367\pm 0.001$ $-38.400\pm 0.002$ $-442.217\pm 0.004$
RandomInit $-54.733\pm 0.001$ $-54.408\pm 2.389$ $-56.560\pm 5.151$ $-790.453\pm 63.779$
No Attack $-20.933\pm 0.000$ $-20.133\pm 1.581$ $-27.590\pm 4.011$ $-109.300\pm 7.603$
@@
DOAR $-32.415\pm 0.000$ $-73.233\pm 0.000$ $-108.369\pm 0.000$ $-1500.085\pm 0.000$
RandomInit $-23.700\pm 0.000$ $-62.097\pm 6.977$ $-68.078\pm 5.959$ $-1470.797\pm 9.742$
No Attack $4.620\pm 0.000$ $-63.658\pm 5.829$ $-66.671\pm 5.255$ $-1486.352\pm 10.795$
@@ 8 Acknowledgements
We would like to thank Doug White and Randy Rinehart for useful industry insights.

### v3

@@ 3 Cybersecurity Simulator
At the high level, our cybersecurity simulator is comprised of four components: 1) a (hierarchically) networked set of devices, 2) workloads, 3) exploits, and 4) attack mitigation actions. In particular, the analysis is centered around a network comprised of subnets; each subnet, in turn, consists of a network of devices. As such, the key unit of interest in our simulator is whether or not particular devices are compromised at particular points in time. Some devices may be more valuable (for example, servers), and some compromises of these more severe than others (for example, exfiltration of trade secrets compared to a requirement to reboot the device). Generally, we wish to ask questions of the form “how many devices, on average, have been compromised between times $X$ and YY?”, “what is the likelihood of particular secrets being exposed to the malicious actor?”, “what is the expected reduction in productivity as a result of cyber-attacks?”, and the like, as we experiment with cyber-attack and defense settings.
@@ 4 Simulation-Based Game-Theoretic Model
We now use the simulation model discussed above to formally define a simulation-based game-theoretic model of cybersecurity encounters. First, there are two players: the defender $\beta$ (e.g., network designer, administrator, etc) and the attacker $\alpha$; these are already explicitly referenced in the simulator described in Section 3). Second, the game takes place over a sequence of discrete time steps $t$ starting at $t=0$ and ending at a predefined horizon $T$. At time 0, the attacker has no knowledge of the network structure $G_{0}$ or device configurations $d_{i}$, and we assume some initial proportion $\eta$ are compromised, all others are clean (i.e., compromise state $c_{i0}$ is a zero vector for all other devices, 1 for the fraction $\eta$). Additionally, we assume that the compromised state $c_{it}$ at any time $t$ after a successful compromise is known to the attacker but not the defender. However, we assume that there is a common knowledge distribution (e.g., stochastic model) of the network generation and device configuration, i.e., $G_{0}\sim\mathcal{D}_{G}$ and $\{d_{i}\}_{i\in D_{0}}\sim\mathcal{D}_{d}$. Consequently, our game entails partial observability of game (device) state by both players.
@@ 4.1 Attacker Actions
Each attacker probe action $a_{q,i,j}^{\alpha}$ is associated with a particular probe $q$ which targets device $j$ from device $i$ (i.e., there is a distinct action for each probe $q$ and device pair $i$ and $j$, although of course only a subset of these have a non-zero probability of success, and it is straightforward to prune this set accordingly to only consider probes across existing network edges). Let $A_{\mathit{probe}}^{\alpha}$ be the set of all such (feasible) combinations (note that a probe may also be guaranteed to fail due to configuration mismatch on the target device $j$, but this information may not be available to the attacker). The strategic decision corresponds to selecting a problem $a^{\alpha}\in A_{\mathit{probe}}^{\alpha}$ at a given time step $t$. The side-effect of each probe is purely informational, providing the attacker with some of the details about the states of devices on the network (but not complete details).
@@ 4.2 Defender Actions
Just as in the case of attacker, the defender’s probing actions target particular devices that are not in the defender’s network to obtain their characteristics. Structurally, these mirror the attacker’s probing actions described above, with the same constraints imposed; thus, each probe targets a device $j$ from a device $i$, so that the full set of probes corresponds to the cross product between probe action types (ping, traceroute, etc) and pairs of devices. We let $A_{\mathit{probe}}^{\beta}$ be the set of feasible probes (i.e., requiring connectivity from $i$ to $j$).
@@ 5 DOAR: Double Oracle with Actor Response Ascent
$\mathrm{noop}\;=\;(t_{\mathrm{noop}},\;\emptyset,\;\emptyset,\;0)$
@@
$D^{*}=\{\,d\mid a_{d}\neq\mathrm{noop}\,\},\quad t^{*}\sim\{\,t_{d}\mid d\in D^{*}\,\},\quad e^{*}=\bigcup_{d}e_{d},\quad p^{*}\sim\{\,p_{d}\mid d\in D^{*}\,\}.$
@@ 6 Experiments
The environment tracks up to $N_{\max}$ devices, each described by six features—OS code, version number, compromise flag, anomaly score, attacker-knowledge flag, and presence indicator—concatenated (and 1-1-padded) into a full state vector $x\in\mathbb{R}^{6N_{\max}}$. The defender’s observation $o_{d}\in\mathbb{R}^{6N_{\max}}$ masks comprimise information as well as attacker owned and devices not yet added to the network via its evolution procedure. The attacker’s observation $o^{\alpha}\in\mathbb{R}^{4N_{\max}}$ further hides anomaly scores and presence indicators of the defender, drops the attacker-knowledge feature entirely, and masks inactive or unknown devices. The defender’s observations $o^{\beta}$ include anomaly scores and device configurations, but not actual compromise flags.
@@
Defender Clean Clears compromise flag, incurs device busy-time +0.30+0.30 if compromised, else 0.01-0.01
Checkpoint Saves state (of all devices) to disk for future rollback 0.50-0.50
Restore Restores network from last checkpoint (losing intermediate work) 1.00-1.00
Upgrade Increments application version, incurs busy-time 1.00-1.00 (device also becomes busy)
Scan Executes ML-based anomaly detection on recent logs 0.50-0.50
Block / Unblock Toggles network edges to control reachability 0.50-0.50
Pass No action taken 0
Attacker Attack Deterministically exploits any matching vulnerability on each targeted host +1.00+1.00 per compromise; +10.0+10.0 if via domain controller
Probe Discovers previously unknown neighbors via network scanning +0.10+0.10 per successful discovery
Pass No action taken 0
@@
[…]
@@ 6.1 Results
DOAR $14.700\pm 0.001$ $97.367\pm 0.001$ $38.400\pm 0.002$ $442.217\pm 0.004$
RandomInit $54.733\pm 0.001$ $54.408\pm 2.389$ $56.560\pm 5.151$ $790.453\pm 63.779$
No Attack $20.933\pm 0.000$ $20.133\pm 1.581$ $27.590\pm 4.011$ $109.300\pm 7.603$
@@
DOAR $32.415\pm 0.000$ $73.233\pm 0.000$ $108.369\pm 0.000$ $1500.085\pm 0.000$
RandomInit $23.700\pm 0.000$ $62.097\pm 6.977$ $68.078\pm 5.959$ $1470.797\pm 9.742$
No Attack $4.620\pm 0.000$ $3.658\pm 5.829 66$$6.671\pm 5.255$ $1486.352\pm 10.795$
@@ 8 Acknowledgements
We would like to thank Doug White and Randy Rinehart for useful industry insights. This research was partially supported by the NSF (IIS-2214141), ONR (N00014-24-1-2663) and ARO (W911NF-25-1-0059).